OpenStudy (anonymous):

A rectangular prism has length thirteen meters, width six meters, and volume three hundred ninety cubic meters. What is the height of the prism?

OpenStudy (cwrw238):

volume of the prism = l*w*h form an equation: 13* 6 * h = 390 now solve for h

OpenStudy (anonymous):

5 meters

OpenStudy (cwrw238):

390 / 78 = 5 correct